# Ordering the topological order in the fractional quantum Hall effect

### v2

@@ Acknowledgements.
We thank Maissam Barkeshli, Jaume Gomis, Zhaoyu Han, Jainendra Jain, Kristan Jensen, Chao-Ming Jian, Anton Kapustin, Zohar Komargodski, Ethan Lake, Ho Tat Lam, Greg Moore, Salvatore Pace, Shu-Heng Shao, Sahand Seifnashri, G. Sreejith, Cenke Xu, and Carolyn Zhang for discussions. MC is partially supported by NSF grant DMR-2424315. SM is supported by the Laboratory for Physical Sciences. AR is partially supported by the U.S. Department of Energy under Grant No. DE-SC0022021. NS is partially supported by DOE grant DE-SC0009988. TS is partially supported by NSF grant DMR-2206305. AR, NS, and TS are also partially supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Grant No. 651678, AK, Grant No. 651444, NS, and Grant No. 651446, TS).
@@ A.1 $\mathbb{Z}_{N}$ gauge theory or $D(\mathbb{Z}_{N})$
Clearly, this braiding is non-degenerate for all values of $N$, and hence, this theory is modular. This braiding means that each $\mathbb{Z}_{N}^{(1)}$ one-form symmetry is anomaly-free, but there is a mixed anomaly between the two $\mathbb{Z}_{N}$ factors. For even $N$, the theory does not factorize in terms of independent $\mathbb{Z}_{N}$ factors. But for odd $N$ it factorizes as $\mathcal{A}^{N,2}\boxtimes\mathcal{A}^{N,-2}$, which are generated by $em$ and $em^{-1}$.
This theory can also be described using the Chern-Simons Lagrangian [158]:
@@
A.2 Spin$(n)_{1}$ and Spin$(n)_{1}\boxtimes\{1,c\}$
Here, Spin$(n)_{1}$ with $n\text{ mod }{16}$ denotes the TQFT of the level-$1$ Spin$(n)$ Chern-Simons theory. Alternatively, it is the $n$th TQFT in Kitaev’s sixteen-fold way [39]. The corresponding RCFTs are theories of $n$ Majorana fermions summed over their spin structures.
@@ A.2 Spin$(n)_{1}$ and Spin$(n)_{1}\boxtimes\{1,c\}$
$\mathrm{Spin}(n)_{1}$ =$=\{1,\sigma,\psi\}$ (A.4)
@@
From this it is clear that $d_{1}=1=d_{\psi}$ and $d_{\sigma}=\sqrt{2}$. Eight choices of exchange and braiding statistics are consistent with these fusion rules, as classified in App. D in [33] and in [39]:
@@
For $n$ an even integer, these are Abelian TQFTs. They were reviewed extensively by Kitaev [39], and we refer the reader to that reference. They happen to coincide with some of our previous definitions:
$\mathrm{Spin}(n)_{1}=\begin{cases}\mathcal{A}^{4,n/2}&\mbox{if }n=2\text{ mod }{4}\\ \mathcal{A}^{2,n/4}\boxtimes\mathcal{A}^{2,n/4}&\mbox{if }n=4\text{ mod }{8}\\ D(\mathbb{Z}_{2})&\mbox{if }n=0\text{ mod }{16}\,.\end{cases}$ (A.11)
@@ A.4 Pfaffian TOs
Let us review a field theory description of the Moore-Read state $\mathsf{Pf}_{2,1,1}^{f}$. Following [115], we start with ${\rm U}(2)_{2,-4}\times\mathrm{U}(1)_{4q}$ Chern-Simons theory:
@@ Appendix B More general coupling of bosonic TQFT to background $\mathrm{U}(1)$
We start this section by showing that what we did in Sec. 4 was essentially the only way to reproduce the correct response theory. In doing so, we will see the role played by some integers $u_{I}$, which define the coupling of the theory to the background $\mathrm{U}(1)$ gauge field. Later, we will work out some relations that allow us to change these $u_{I}$ without changing the TQFT. This generalizes the discussion about changing the generator of the one-form symmetry in Sec. 7. For simplicity, we will limit ourselves to bosonic theories. The extension to electronic theories is straightforward.
@@
As we vary $A$, the background fields ${\cal B}_{I}$ vary and select a $\mathbb{Z}_{s}^{(1)}$ subgroup of the one-form symmetry group $\otimes_{I}\mathbb{Z}_{s_{I}}^{(1)}$. Specifically, let $X_{I}$ be the generator of $\mathbb{Z}_{s_{I}}$. Then, the fact that all ${\cal B}_{I}$ are determined by the same $dA$, means that the group elements that $A$ couples to are powers of $X=\prod_{I}X_{I}^{u_{I}}$, which satisfies $X^{s}=1$ with $s$ the smallest nonzero solution of $su_{I}=0\text{ mod }s_{I}$ for all $I$. Furthermore, the background $\cal B$ that couples to this $X$ satisfies, as in (4.6), ${\cal B}=\left[dA/2\pi\right]_{s}$.
@@
In more physical terms, we recognize $X$ as associated with the vison $v\in{\cal A}$. Then, the anomaly expression for general ${\cal B}_{I}$ determines $r$ for $\cal B$ and therefore $\sigma_{H}=r/s$. The vison generates the Abelian set of anyons ${\cal V}^{s,r}$ discussed in Sec. 7.
@@
It generates a $\mathbb{Z}_{s^{\prime}}^{(1)}\subseteq\mathbb{Z}_{s}^{(1)}$ with $s^{\prime}={s/\gcd(s,u)}$ one-form symmetry with anomaly $r^{\prime}={ru^{2}/\gcd(s,u)}$ and $u$-parameter $1$, i.e., ${\cal V}^{s^{\prime},r^{\prime}}$. We see that
@@
Let us demonstrate it in a simple example with $\sigma_{H}={1/2}$. We need to have ${r^{\prime}/s^{\prime}}={ru^{2}/s}={1/2}$. The simplest option is ${\cal A}^{2,1}_{1}\cong{\cal V}^{2,1}$. This is $\mathrm{U}(1)_{2}$ with the vison $v=a_{1}$ with $h={1/4}$ and $Q={1/2}$. Alternatively, we can have ${\cal A}^{8,1}_{2}$, i.e., $\mathrm{U}(1)_{8}$ with the vison $v=a_{1}^{2}$ also with $h={1/4}$ and $Q={1/2}$. Either way, the vison generates ${\cal V}^{4,2}$.
@@ B.2 Identifications of ${\cal A}^{s,r}_{u}$
The first relation is the same as in the bosonic theory before we coupled to $A$ (7.9). The second relation reflects the fact that $u$ labels the vison and our labeling of anyons is by $j\sim j+s$ [^45] We could have replaced the second identification with a stronger version $\mathcal{A}^{s,r}_{u}\cong\mathcal{A}^{s,r}_{u+{s/\ell}}$. We did not do it because even though this does not change $h$ and $Q$ of the anyons in $\mathcal{A}^{s,r}_{u}$, it does change the values of $Q$ of other anyons in $\cal T$.
@@ C.1 Necessary definitions and results
MTCs describe line operators in a bosonic TQFT. Super-modular categories describe line operators in a spin TQFT. In this case, it may appear that one should identify $1$ and $c$, and reduce the super-modular category to an even smaller theory with half the number of anyons. For example, the Moore-Read state can be described by a super-modular category with 12 anyon types (including $c$). After identifying $c$ and $1$, one can write down a theory of 6 anyons. However, it is known that this theory does not admit any consistent braiding [98], i.e., it is merely a unitary fusion category, but not braided. Thus, we will often keep $c$ as an element of ${\cal T}$ throughout this appendix.
@@
A super-modular category splits if it can be written as ${\cal M}\boxtimes\{1,c\}$, where $\cal M$ is an MTC.
@@
It will be crucial for our later discussion to consider the minimal modular extension ${\cal T}^{\rm ext}$ of a super-modular category $\cal T$. The minimal modular extension is a minimal (by total quantum dimension) MTC $\cal{T}^{\rm ext}$ that contains $\cal{T}$ as a subcategory. It is $\mathbb{Z}_{2}$ graded: ${\cal T}^{\rm ext}={\cal T}_{0}^{\rm ext}\oplus{\cal T}_{1}^{\rm ext}$, where ${\cal T}_{0}^{\rm ext}$ contains all anyons braiding trivially with $c$, and ${\cal T}_{1}^{\rm ext}$ contains all anyons which have $-1$ braiding with $c$. Moreover, ${\cal T}_{0}^{\rm ext}={\cal T}$. Then using Lemma D.2, the $\mathbb{Z}_{2}$ grading implies that $\mathcal{D}_{\cal{T}^{\rm ext}}=\sqrt{2}\mathcal{D}_{\cal{T}}$. Mathematically, the existence of a minimal modular extension for any super-modular category was recently established in [162].
We will sometimes refer to anyons in $\mathcal{T}^{\rm ext}_{1}$ as “Ramond anyons/lines.” [^47] These lines correspond to states in the Hilbert space of the theory on a 2d torus with spin structures (+,+)(+,+)( + , + ) and (+,)(+,-)( + , - ). Here, the two signs represent the boundary conditions of spinors around the two cycles of the torus. $+$ means that they are periodic (Ramond or R in the string theory terminology), and -- means that they are anti-periodic (Neveu–Schwarz or NS in the string theory terminology). The first sign corresponds to the boundary conditions as the spinors circle around the line, and the second sign corresponds to the boundary conditions along the line. We will denote them by bold fonts, e.g., $\bm{a}$. Ramond anyons come in two types: Majorana and non-Majorana. An anyon $\bm{a}\in\mathcal{T}^{\rm ext}_{1}$ is Majorana if $c\times\bm{a}=\bm{a}$, otherwise it is non-Majorana. The Majorana anyons correspond to states with odd fermion parity on the torus in the (+,+)(+,+)( + , + ) sector, which will be discussed further in App. C.2.
@@
Without gauging/summing over the spin structure, the Ramond “anyons” are not actual anyonic excitations in the fermionic topological order because they have $-1$ braiding with the fermions. In other words, they are extrinsic objects carrying $\pi$ fluxes (in the sense of Aharonov-Bohm phase). A concrete example is the vortices in an (electronic) superconductor. In the TQFT, they are line operators attached to fermion parity surfaces. After gauging, they become genuine line operators of the bosonic shadow. [^48] Here we follow the terminology of [69], according to which non-genuine lines are lines attached to topological surfaces. The non-genuine lines, like this Ramond line, are higher-dimensional versions of the familiar disorder operator in 1+1d. The latter is a non-genuine point operator, because it is attached to a topological line, which is the symmetry line of a global $\mathbb{Z}_{2}$ symmetry.
@@
This can be understood as stacking $n$ copies of $p+ip$ superconductors before gauging. Equivalently, as a TQFT, different choices of the modular extension differ by phase factors in the sum over spin structures (see App. C.5 in [115]). [^49] In various contexts, the freedom in such phases in the process of gauging is known as discrete torsion or as adding Dijkgraaf-Witten terms. Equivalently, they correspond to adding local counterterms or tensoring an invertible field theory before the gauging.
@@ C.2 The Hilbert space of the TQFT on a two-torus
Consider the TQFT on a three-torus. It has eight spin structures labeled in an obvious way as (±,±,±)plus-or-minusplus-or-minusplus-or-minus($\pm$,$\pm$,$\pm$)( ± , ± , ± ). The (+,+,+)(+,+,+)( + , + , + ) spin structure is referred to as odd, and the other seven spin structures are even. We can change the cycles of the torus using SL$(3,\mathbb{Z})$ transformations. This has the effect of permuting the even spin structures, while leaving the odd spin structure invariant.
In a spin TQFT, the partition function depends on the spin structure. We’ll denote it as ${\cal Z}_{(\pm,\pm,\pm)}$. Using SL$(3,\mathbb{Z})$ transformations ${\cal Z}_{(\pm,\pm,\pm)}$ can have only two different values. The partition functions for the even spin structures are $N_{e}$ and ${\cal Z}_{(+,+,+)}=N_{o}$.
These partition functions can be interpreted as traces over the Hilbert space of a spatial two-torus. The four spin structures of the spatial torus are labeled by (±,±)plus-or-minusplus-or-minus($\pm$,$\pm$)( ± , ± ), and their Hilbert spaces are ${\cal H}_{(\pm,\pm)}$. Interpreting the first index in ${\cal Z}_{(\pm,\pm,\pm)}$ as the spin structure around the compact Euclidean time and the other ones as associated with the spatial cycles, we have
$\begin{split}&{\cal Z}_{(-,\pm,\pm)}={\rm Tr\,}_{{\cal H}_{(\pm,\pm)}}1=N_{e}\\ &{\cal Z}_{(+,-,+)}={\rm Tr\,}_{{\cal H}_{(-,+)}}(-1)^{F}=N_{e}\\ &{\cal Z}_{(+,+,-)}={\rm Tr\,}_{{\cal H}_{(+,-)}}(-1)^{F}=N_{e}\\ &{\cal Z}_{(+,-,-)}={\rm Tr\,}_{{\cal H}_{(-,-)}}(-1)^{F}=N_{e}\\ &{\cal Z}_{(+,+,+)}={\rm Tr\,}_{{\cal H_{(+,+)}}}(-1)^{F}=N_{o}\,,\end{split}$ (C.5)
where $(-1)^{F}$ is the fermion parity. We immediately learn some facts about these four Hilbert spaces ${\cal H}_{(\pm,\pm)}$:
  • The dimensions of ${\cal H}_{(\pm,\pm)}$ are the same and are given by the positive integer $N_{e}$.
@@
  • The Hilbert space of the odd spatial spin structure ${\cal H}_{(+,+)}$ has $N_{e}+N_{o}\over 2$ states with $(-1)^{F}=+1$ and $N_{e}-N_{o}\over 2$ states with $(-1)^{F}=-1$. Therefore, $N_{o}$ must be an integer satisfying $N_{o}=N_{e}\text{ mod }2$ and $|N_{o}|\leq N_{e}$.
These integers are related to the data of the bosonic shadow [165]. (See App. C.1.) Denote the number of anyon types (including $c$) in the super-modular category $\mathcal{T}$ by $|\mathcal{T}|$, and similarly define $|\mathcal{T}^{\rm ext}_{0,1}|$. Further we denote the number of Majorana-type anyons in $\mathcal{T}^{\rm ext}_{1}$ as $|\mathcal{T}^{\rm ext}_{\rm M}|$. Then we have
@@
(Note that since $\mathcal{T}$ is a super-modular category and it includes $c$, $|\mathcal{T}|$ must be even.) In other words, the number of $(-1)^{F}=-1$ states in ${\cal H}_{(+,+)}$ is $|\mathcal{T}^{\rm ext}_{\rm M}|$.
One way to understand this is the following. Observe that $(-1)^{F}=-1$ means that there is a single $c$ defect on the torus. Therefore, such a state is obtained by inserting a defect of Majorana-type anyon in the bosonic shadow inside the solid torus and performing the path integral. This explains why the number of such states is equal to $|\mathcal{T}^{\rm ext}_{\rm M}|$. It is worth noting that the TQFT states in the (+,+)(+,+)( + , + ) sector depend to some extent on the choice of the minimal modular extension, or equivalently the chiral central charge (namely, for modular extensions given in (C.4), $N_{o}$ depends only on the even/oddness of $n$).
If the TQFT is also a spinc theory, the previous discussion is still valid, but we can also turn on background $A$. That background allows us to interpolate continuously between the distinct spin structures without $A$ we discussed above. Let us turn on only spatial $A$. The expressions above in terms of traces are still valid. And since they are given by integers, they cannot vary with $A$. Therefore, $N_{e}=N_{o}$. This means that none of the four Hilbert spaces ${\cal H}_{(\pm,\pm)}$ have any states with $(-1)^{F}=-1$. In particular, it implies that there are no Majorana anyons. A simple way to understand it follows from the definition of Majorana anyons as anyons $\bm{a}$, such that $c\times\bm{a}=\bm{a}$. In a spinc theory, $c$ carries charge one under $\mathrm{U}(1)$ and therefore there cannot be any such $\bm{a}$. We will discuss it in more detail in Sec. b.
For example, $\mathrm{U}(1)_{k}$ with odd $k$ is a spinc theory with $N_{e}=N_{o}=k$. It has $k$ states on a spatial two-torus for each spin structure, all with $(-1)^{F}=+1$ [165]. On the other hand, ${\rm SO}(3)_{k}={{\rm SU}(2)_{k}\over\mathbb{Z}_{2}}$ with $k=2\text{ mod }4$ is a spin theory, but it is not a spinc theory. It has $N_{e}={k+2\over 4}$ and $N_{o}={k-6\over 4}$. This corresponds to ${k+2\over 4}$ states for each spin structure on a spatial two-torus and only one state with $(-1)^{F}=-1$ in the (+,+)(+,+)( + , + ) spin structure [165].
The key fact we used about spinc theories is that, unlike ordinary (non-spinc) fermionic theories, in these cases, we can interpolate continuously between the different spin structures. This continuous interpolation is a generalization of the flux threading argument, reviewed in Sec. 2.2. The original $2\pi$-flux threading maps the theory back to itself, perhaps with some nontrivial action on the Hilbert space. In spinc theories, we can extend it by threading $\pi$-flux. It does not map the theory back to itself, but it changes the spatial spin structure, thus relating the Hilbert spaces with different spin structures.

### v3

@@ Acknowledgements.
We thank Maissam Barkeshli, Jaume Gomis, Zhaoyu Han, Jainendra Jain, Kristan Jensen, Chao-Ming Jian, Anton Kapustin, Zohar Komargodski, Ethan Lake, Ho Tat Lam, Greg Moore, Salvatore Pace, Shu-Heng Shao, Sahand Seifnashri, G. Sreejith, Cenke Xu, and Carolyn Zhang for discussions. MC is partially supported by NSF Grant DMR-2424315. SM is supported by the Laboratory for Physical Sciences. AR is partially supported by the U.S. Department of Energy under Grant No. DE-SC0022021. NS is partially supported by DOE Grant DE-SC0009988. TS is partially supported by NSF Grant DMR-2206305. AR, NS, and TS are also partially supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Grant No. 651678, AK; Grant No. 651444, NS; and Grant No. 651446, TS).
@@ A.1 $\mathbb{Z}_{N}$ gauge theory or $D(\mathbb{Z}_{N})$
Clearly, this braiding is nondegenerate for all values of $N$, and hence, this theory is modular. This braiding means that each $\mathbb{Z}_{N}^{(1)}$ one-form symmetry is anomaly free, but there is a mixed anomaly between the two $\mathbb{Z}_{N}$ factors. For even $N$, the theory does not factorize in terms of independent $\mathbb{Z}_{N}$ factors. But for odd $N$ it factorizes as $\mathcal{A}^{N,2}\boxtimes\mathcal{A}^{N,-2}$, which are generated by $em$ and $em^{-1}$.
This theory can also be described using the Chern-Simons Lagrangian [156],
@@
[…]
@@
$\mathrm{Spin}(n)_{1}$ $=\{1,\sigma,\psi\}$ (A.4)
@@
From this it is clear that $d_{1}=1=d_{\psi}$ and $d_{\sigma}=\sqrt{2}$. Eight choices of exchange and braiding statistics are consistent with these fusion rules, as classified in Appendix D in [35] and in [41],
@@
For $n$ an even integer, these are Abelian TQFTs. They were reviewed extensively by Kitaev [41], and we refer the reader to that reference. They happen to coincide with some of our previous definitions,
[…]
@@ A.4 Pfaffian TOs
Let us review a field theory description of the Moore-Read state $\mathsf{Pf}_{2,1,1}^{f}$. Following [114], we start with ${\rm U}(2)_{2,-4}\times\mathrm{U}(1)_{4q}$ Chern-Simons theory,
@@ Appendix B More general coupling of bosonic TQFT to background $\mathrm{U}(1)$
We start this appendix by showing that what we did in Sec. 4 was essentially the only way to reproduce the correct response theory. In doing so, we will see the role played by some integers $u_{I}$, which define the coupling of the theory to the background $\mathrm{U}(1)$ gauge field. Later, we will work out some relations that allow us to change these $u_{I}$ without changing the TQFT. This generalizes the discussion about changing the generator of the one-form symmetry in Sec. 7. For simplicity, we will limit ourselves to bosonic theories. The extension to electronic theories is straightforward.
@@
[…]
@@
[…]
@@
It generates a $\mathbb{Z}_{s^{\prime}}^{(1)}\subseteq\mathbb{Z}_{s}^{(1)}$ with $s^{\prime}={s/\gcd(s,u)}$ one-form symmetry with anomaly $r^{\prime}={ru^{2}/\gcd(s,u)}$ and $u$-parameter 11, i.e., ${\cal V}^{s^{\prime},r^{\prime}}$. We see that
@@
Let us demonstrate it in a simple example with $\sigma_{H}={1/2}$. We need to have ${r^{\prime}/s^{\prime}}={ru^{2}/s}={1/2}$. The simplest option is ${\cal A}^{2,1}_{1}\cong{\cal V}^{2,1}$. This is $\mathrm{U}(1)_{2}$ with the vison $v=a_{1}$ with h=1/4h={1/4} and Q=1/2Q={1/2}. Alternatively, we can have ${\cal A}^{8,1}_{2}$, i.e., $\mathrm{U}(1)_{8}$ with the vison $v=a_{1}^{2}$ also with h=1/4h={1/4} and Q=1/2Q={1/2}. Either way, the vison generates ${\cal V}^{4,2}$.
@@ B.2 Identifications of ${\cal A}^{s,r}_{u}$
The first relation is the same as in the bosonic theory before we coupled to $A$ (7.9). The second relation reflects the fact that $u$ labels the vison and our labeling of anyons is by $j\sim j+s$. [^45] We could have replaced the second identification with a stronger version $\mathcal{A}^{s,r}_{u}\cong\mathcal{A}^{s,r}_{u+{s/\ell}}$. We did not do it because even though this does not change $h$ and $Q$ of the anyons in $\mathcal{A}^{s,r}_{u}$, it does change the values of $Q$ of other anyons in $\cal T$.
@@ C.1 Necessary definitions and results
MTCs describe line operators in a bosonic TQFT. Super-modular categories describe line operators in a spin TQFT. In this case, it may appear that one should identify $1$ and $c$, and reduce the super-modular category to an even smaller theory with half the number of anyons. For example, the Moore-Read state can be described by a super-modular category with 12 anyon types (including $c$). After identifying $c$ and $1$, one can write down a theory of six anyons. However, it is known that this theory does not admit any consistent braiding [100], i.e., it is merely a unitary fusion category, but not braided. Thus, we will often keep $c$ as an element of $\cal T$} throughout this appendix.
@@
[…]
@@
[…]
We will sometimes refer to anyons in $\mathcal{T}^{\rm ext}_{1}$ as “Ramond anyons/lines.” [^47] These lines correspond to states in the Hilbert space of the theory on a 2D torus with spin structures $(+,+)$ and (+,)(+,-). Here, the two signs represent the boundary conditions of spinors around the two cycles of the torus. ++ means that they are periodic (Ramond or R in the string theory terminology), and - means that they are anti-periodic (Neveu–Schwarz or NS in the string theory terminology). The first sign corresponds to the boundary conditions as the spinors circle around the line, and the second sign corresponds to the boundary conditions along the line. We will denote them by bold fonts, e.g., $\bm{a}$. Ramond anyons come in two types: Majorana and non-Majorana. An anyon $\bm{a}\in\mathcal{T}^{\rm ext}_{1}$ is Majorana if $c\times\bm{a}=\bm{a}$, otherwise it is non-Majorana. The Majorana anyons correspond to states with odd fermion parity on the torus in the $(+,+)$ sector, which will be discussed further in Appendix C.2.
@@
Without gauging/summing over the spin structure, the Ramond “anyons” are not actual anyonic excitations in the fermionic topological order because they have 1-1 braiding with the fermions. In other words, they are extrinsic objects carrying $\pi$ fluxes (in the sense of Aharonov-Bohm phase). A concrete example is the vortices in an (electronic) superconductor. In the TQFT, they are line operators attached to fermion parity surfaces. After gauging, they become genuine line operators of the bosonic shadow. [^48] Here we follow the terminology of [71], according to which nongenuine lines are lines attached to topological surfaces. The nongenuine lines, like this Ramond line, are higher-dimensional versions of the familiar disorder operator in 1+1d. The latter is a non-genuine point operator, because it is attached to a topological line, which is the symmetry line of a global $\mathbb{Z}_{2}$ symmetry.
@@
This can be understood as stacking $n$ copies of $p+ip$ superconductors before gauging. Equivalently, as a TQFT, different choices of the modular extension differ by phase factors in the sum over spin structures (see Appendix C.5 in [114]). [^49] In various contexts, the freedom in such phases in the process of gauging is known as discrete torsion or as adding Dijkgraaf-Witten terms. Equivalently, they correspond to adding local counterterms or tensoring an invertible field theory before the gauging.
@@ C.2 The Hilbert space of the TQFT on a two-torus
[…]
@@
[…]
These integers are related to the data of the bosonic shadow [163]. (See Appendix C.1.) Denote the number of anyon types (including $c$) in the super-modular category $\mathcal{T}$ by $|\mathcal{T}|$, and similarly define $|\mathcal{T}^{\rm ext}_{0,1}|$. Further we denote the number of Majorana-type anyons in $\mathcal{T}^{\rm ext}_{1}$ as $|\mathcal{T}^{\rm ext}_{\rm M}|$. Then we have
@@
[…]
One way to understand this is the following. Observe that $(-1)^{F}=-1$ means that there is a single $c$ defect on the torus. Therefore, such a state is obtained by inserting a defect of Majorana-type anyon in the bosonic shadow inside the solid torus and performing the path integral. This explains why the number of such states is equal to $|\mathcal{T}^{\rm ext}_{\rm M}|$. It is worth noting that the TQFT states in the $(+,+)$ sector depend to some extent on the choice of the minimal modular extension, or equivalently the chiral central charge [namely, for modular extensions given in (C.4), $N_{o}$ depends only on the even/oddness of $n$].
[…]
For example, $\mathrm{U}(1)_{k}$ with odd $k$ is a spinc theory with $N_{e}=N_{o}=k$. It has $k$ states on a spatial two-torus for each spin structure, all with $(-1)^{F}=+1$ [163]. On the other hand, ${\rm SO}(3)_{k}={{\rm SU}(2)_{k}\over\mathbb{Z}_{2}}$ with $k=2\text{ mod }4$ is a spin theory, but it is not a spinc theory. It has $N_{e}={k+2\over 4}$ and $N_{o}={k-6\over 4}$. This corresponds to $k+2\over 4$} states for each spin structure on a spatial two-torus and only one state with $(-1)^{F}=-1$ in the $(+,+)$ spin structure [163].
The key fact we used about spinc theories is that, unlike ordinary (non-spinc) fermionic theories, in these cases, we can interpolate continuously between the different spin structures. This continuous interpolation is a generalization of the flux threading argument, reviewed in Sec. 2.2. The original $2\pi$-flux threading maps the theory back to itself, perhaps with some nontrivial action on the Hilbert space. In spinc theories, we can extend it by threading $\pi$ flux. It does not map the theory back to itself, but it changes the spatial spin structure, thus relating the Hilbert spaces with different spin structures.